OpenStudy (anonymous):

What is the value of y? 6(5 + 3) = 6 • 5 + 6 • y

OpenStudy (anonymous):

i get 1 1/3 but is there another answer that i can not find???

OpenStudy (anonymous):

@Peaches15 @nincompoop

OpenStudy (anonymous):

if anyone dose not know then do you now anyone who can help???

OpenStudy (anonymous):

@Michele_Laino

OpenStudy (michele_laino):

here we can to compute the left side as below: 6(5+3)=6*5+6*3=...?

OpenStudy (michele_laino):

I got a different solution.

OpenStudy (anonymous):

what would that be and just to let you now my computer is acting up on me

OpenStudy (michele_laino):

hint: if I simplify the left side, I get: \[\Large 30 + 18 = 30 + 6y\] now I subtract at both sides 30, so I can write this: \[\Large 30 + 18 - 30 = 30 + 6y - 30\] please simplify

OpenStudy (mathstudent55):

This problem is an exercise on the distributive property. \(\Large a(b + c) = ab + ac\) Here is an example with numbers: \(\Large 3(4 + 7) = 3 \cdot 4 + 3 \cdot 7\) You take the number outside the parentheses, and you multiply it by both numbers inside the parentheses. You separate the two multiplications by the same operation that is inside the parentheses.

OpenStudy (anonymous):

i so far got 18=6y is that correct @Michele_Laino

OpenStudy (mathstudent55):

Now let's look at your problem: \(\Large 6(5 + 3)\) Let's apply the distributive property to it: \(\Large 6(5 + 3) = 6 \cdot 5 + 6 \cdot \_\_\) What goes in the blank?

OpenStudy (anonymous):

so 3=y

OpenStudy (michele_laino):

yes! now, please I divide both sides by 6, so I get: \[\Large \frac{{18}}{6} = \frac{{6y}}{6}\] again, please simplify

OpenStudy (anonymous):

so 3=y

OpenStudy (michele_laino):

correct!

OpenStudy (anonymous):

thanks

OpenStudy (michele_laino):

:)